### v1

@@ 4.2. A remark on maximal coarse Baum-Connes conjecture
In this section, we discuss how the method used in [HIT2020] to prove the coarse Baum-Connes conjecture also applies to the maximal coarse Baum-Connes conjecture. This requires the following proposition.
@@
Notice that $\ell^{\infty}(X,\mathcal{K})\subseteq\mathbb{C}[X]$ is a subalgebra. Moreover $\ell^{\infty}(X,\mathcal{B})$ is identified with the multiplier algebra of $\ell^{\infty}(X,\mathcal{K})$. Any *-representation of a $C^{*}$-algebra $A$ can be lift to a representation of $M(A)$, see [HIT2020, Proposition 1.7.3]. We shall denote $\rho:\ell^{\infty}(X,\mathcal{B})\to\mathcal{B}(H)$ the lifted representation of $\pi|_{\ell^{\infty}(X,\mathcal{K})}$.
By [GWY2008, Lemma 3.4], any element $T\in\mathbb{C}[X]$ can be written in the form
@@
By the above proposition, we can apply the methods from [HIT2020, Chapter 12] and [WXYZ2024] to the maximal coarse Baum-Connes conjecture. First, consider a maximal representation $\pi$ of the Roe algebra $\mathbb{C}[X,\mathcal{H}_{X}]$. By 4.5, since $\mathbb{C}[X,\mathcal{H}_{X}]$ and $\mathbb{C}[X,\mathcal{H}_{X,E}]$ are unitarily equivalent, this induces a maximal representation $\pi_{E}$ of $\mathbb{C}[X,\mathcal{H}_{X,E}]$. One can identify both $\ell^{\infty}(X)$ and $C_{0}(E)$ as an subalgebra of $\ell^{\infty}(X,\mathcal{B}(L^{2}(E,\mathfrak{h}_{E})\otimes\mathcal{H}))$. Therefore, even in the maximal Roe algebra, we can still define the propagation with respect to $X$ and $E$. By 4.6, we can write an element $T\in C^{*}_{\max}(X,\mathcal{H}_{X,E})$ as a sequence of $T_{n}$ with each $T_{n}=\chi_{X_{n}}T\chi_{X_{n}}$. Moreover, both $\chi_{R,p,v}$ in 3.4 and $F_{s,n}$ in 3.5 are elements in $\ell^{\infty}(X,\mathcal{B}(L^{2}(E,\mathfrak{h}_{E})\otimes\mathcal{H}))$, the condition (2) in 3.4 remains well-defined for the maximal Roe algebra.
Recalling the proof of 3.6 in [HIT2020, Chapter 12], one finds that the proofs of these statements are, in fact, carried out before completion, namely, within the algebras $\mathbb{C}[X,\mathcal{H}_{X}]$ and $\mathbb{A}[X,E]$. For example, $F$ is a multiplier of $\mathbb{A}[X,E]$, $\mathbb{C}[X,\mathcal{H}_{X}]$ is a subalgebra of the multiplier algebra of $\mathbb{A}[X,E]$, and so on. These statements therefore extend to the maximal completions by 4.5. Consequently, we can still define the index map from the maximal Roe algebra to the maximal version of the twisted algebra. Moreover, one can verify that the proofs in [HIT2020, Section 12.3 & 12.4] also hold within the dense subalgebra $\mathbb{C}[X,\mathcal{H}_{X,E}]$, and thus they naturally extend to the maximal completions. Hence, combining these observations, we obtain:
@@
    The canonical quotient map $\pi:C^{*}_{\max}(X)\to C^{*}(X)$ induces an isomorphism on $K$-theory, i.e.,
@@
    For any subspace $Y$, the canonical quotient map $\pi:C^{*}_{\max,Y,\infty}(X)\to C^{*}_{Y,\infty}(X)$ induces an isomorphism on the level of $K$-theory, i.e.,
@@
    For any subspace $Y$, the maximal relative coarse Baum-Connes conjecture holds for $(X,Y)$.
@@
by taking quotient pointwisely. Since both maximal and reduce coarse Baum-Connes conjecture holds for $X$ (by 4.7 and [WXYZ2024, Theorem 1.1]), the maps $\mu_{\max}$ and $\mu$ are isomorphisms. Thus, it suffices to show that $\pi_{L}$ induces an isomorphism on $K$-theory, this holds from a cutting and pasting argument in [Yu1997, Theorem 3.2].
@@
where $C^{*}_{\max}(X,Y)$ denotes the geometric ideal in the maximal Roe algebra. Its $K$-theory is isomorphic to $K_{*}(C^{*}_{\max}(Y))$; see [GWZ2025] for the associated short exact sequence. For the proof of this isomorphism, one may refer to [HRY1993, Lemma 1] for the reduced case. The argument for the maximal case follows verbatim.
The left vertical map is given by the compostion

### v2

@@ 4.2. A remark on maximal coarse Baum-Connes conjecture
In this section, we discuss how the method used in Willett and Yu [2020] to prove the coarse Baum-Connes conjecture also applies to the maximal coarse Baum-Connes conjecture. This requires the following proposition.
@@
Notice that $\ell^{\infty}(X,\mathcal{K})\subseteq\mathbb{C}[X]$ is a subalgebra. Moreover $\ell^{\infty}(X,\mathcal{B})$ is identified with the multiplier algebra of $\ell^{\infty}(X,\mathcal{K})$. Any *-representation of a $C^{*}$-algebra $A$ can be lift to a representation of $M(A)$, see [Willett and Yu, 2020, Proposition 1.7.3]. We shall denote $\rho:\ell^{\infty}(X,\mathcal{B})\to\mathcal{B}(H)$ the lifted representation of $\pi|_{\ell^{\infty}(X,\mathcal{K})}$.
By [Gong et al., 2008, Lemma 3.4], any element $T\in\mathbb{C}[X]$ can be written in the form
@@
By the above proposition, we can apply the methods from [Willett and Yu, 2020, Chapter 12] and Wang et al. [2024] to the maximal coarse Baum-Connes conjecture. First, consider a maximal representation $\pi$ of the Roe algebra $\mathbb{C}[X,\mathcal{H}_{X}]$. By 4.5, since $\mathbb{C}[X,\mathcal{H}_{X}]$ and $\mathbb{C}[X,\mathcal{H}_{X,E}]$ are unitarily equivalent, this induces a maximal representation $\pi_{E}$ of $\mathbb{C}[X,\mathcal{H}_{X,E}]$. One can identify both $\ell^{\infty}(X)$ and $C_{0}(E)$ as an subalgebra of $\ell^{\infty}(X,\mathcal{B}(L^{2}(E,\mathfrak{h}_{E})\otimes\mathcal{H}))$. Therefore, even in the maximal Roe algebra, we can still define the propagation with respect to $X$ and $E$. By 4.6, we can write an element $T\in C^{*}_{\max}(X,\mathcal{H}_{X,E})$ as a sequence of $T_{n}$ with each $T_{n}=\chi_{X_{n}}T\chi_{X_{n}}$. Moreover, both $\chi_{R,p,v}$ in 3.5 and $F_{s,n}$ in 3.6 are elements in $\ell^{\infty}(X,\mathcal{B}(L^{2}(E,\mathfrak{h}_{E})\otimes\mathcal{H}))$, the condition (2) in 3.5 remains well-defined for the maximal Roe algebra.
Recalling the proof of 3.7 in [Willett and Yu, 2020, Chapter 12], one finds that the proofs of these statements are, in fact, carried out before completion, namely, within the algebras $\mathbb{C}[X,\mathcal{H}_{X}]$ and $\mathbb{A}[X,E]$. For example, $F$ is a multiplier of $\mathbb{A}[X,E]$, $\mathbb{C}[X,\mathcal{H}_{X}]$ is a subalgebra of the multiplier algebra of $\mathbb{A}[X,E]$, and so on. These statements, therefore, extend to the maximal completions by 4.5. Consequently, we can still define the index map from the maximal Roe algebra to the maximal version of the twisted algebra. Moreover, one can verify that the proofs in [Willett and Yu, 2020, Section 12.3 & 12.4] also hold within the dense subalgebra $\mathbb{C}[X,\mathcal{H}_{X,E}]$, and thus they naturally extend to the maximal completions. Hence, combining these observations, we obtain:
@@
    The canonical quotient map $\pi:C^{*}_{\max}(X)\to C^{*}(X)$ induces an isomorphism in $K$-theory, i.e.,
@@
    For any subspace $Y$ in $X$, the canonical quotient map $\pi:C^{*}_{\max,Y,\infty}(X)\to C^{*}_{Y,\infty}(X)$ induces an isomorphism in $K$-theory, i.e.,
@@
    For any subspace $Y$ in $X$, the maximal relative coarse Baum-Connes conjecture holds for $(X,Y)$.
@@
by taking the quotient pointwise. Since both maximal and reduce coarse Baum-Connes conjecture holds for $X$ (by 4.7 and [Wang et al., 2024, Theorem 1.1]), the maps $\mu_{\max}$ and $\mu$ are isomorphisms. Thus, it suffices to show that $\pi_{L}$ induces an isomorphism in $K$-theory, this holds from a cutting and pasting argument in [Yu, 1997, Theorem 3.2].
@@
where $C^{*}_{\max}(X,Y)$ denotes the geometric ideal in the maximal Roe algebra. Its $K$-theory is isomorphic to $K_{*}(C^{*}_{\max}(Y))$; see Guo et al. [2025] for the associated short exact sequence. For the proof of this isomorphism, one may refer to [Higson et al., 1993, Lemma 1] for the reduced case. The argument for the maximal case follows verbatim.
The left vertical map is given by the composition
@@
Acknowledgement
We thank Jiawen Zhang for insightful comments on the first draft.
Conflict of interest
The authors declare that they have no conflict of interest.
Data availability
Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.